# How typical is contextuality?

### v1

@@ 1.2 Methods
To study the typicality of quantum contextuality, we therefore perform extensive simulations of contextuality certification. We characterize each simulation by a tuple of parameters $(n,m,d;N)$, where $n$ is the number of preparations and $m$ is the number of binary measurements that comprise the prepare-and-measure scenario, $d$ is the dimension of the Hilbert space associated to the quantum system being investigated, and $N$ is the number of sampling iterations and assessments of contextuality. We begin by randomly sampling $n$ density matrices of dimension $d\times d$ and $m$ new density matrices, which together with their complements constitute the $m$ binary measurements of the scenario [^2] As we elaborate on in Ap. B, this sampling method for effects actually samples normalized POVM elements. This is not necessarily the most general sampling one could perform, and we will include results with more generic sampling methods in a future version of this manuscript. Note, however, that this does not change assessments of nonclassicality by the linear program, as nonclassicality depends only on the cone generated by such effects, which is the same if the effects were not normalized (but still non-zero)selby23fragments . (amounting to a total of $2m$ effects). We then employ the linear program selby2024linear to determine whether the statistics specified by $(n,m,d)$ admit a generalized-noncontextual model.
@@
Notice that, throughout this work, stating $t_{(n,m,d;N)}\approx 100\%$ does not mean that contextuality will always be found for $n$ preparations and $m$ binary measurements for a quantum system of dimension $d$, but rather that out of $N$ tries, every single one of them exhibited contextuality by our certification criteria. To calculate the confidence level of the true value of $t_{(n,m,d)}$, we use the Wilson score interval wilson , which we explain in detail in Appendix C. In summary, when we say $t_{(n,m,d;N=10^{6})}\approx 100\%$ in this work, this criterion ensures that the typicality of contextuality $t_{(n,m,d)}$ is at least 99.999% with 99% level of confidence. In this paper, we focus on qubit states $(d=2)$, but the general approach and the simulation toolbox that we share in an online repository are also suitable for higher dimensions [^3] To apply to higher-dimensional systems or to many-outcome measurements, the scripts for sampling randomly would need to be extended. This will be provided in a future version of the repository..
To generate random states and effects, we employ the qutip toolbox lambert2024qutip5quantumtoolbox . Pure random states distributed with a Haar measure are generated by applying a Haar random unitary to a fixed pure state (rand_ket function). To generate mixed states, we follow the algorithm designed in Ref. Bruzda2009 that generates Ginibre random density operators of fixed rank (rand_dm function). Since only full-rank density matrices have nonzero measure Bengtsson2017 , sampling from mixed states will never lead to pure states. More details on random sampling are provided in Appendix B. To sample the effects, there are two approaches that we employ depending on the task— randomly sampling effects or fixing a large number of (somewhat uniform) effects. For the first approach, we sample the effects with the same function that samples the density operators, which yields a unit-trace POVM element $M$. For each sampled operator $M$, we also include its complement $\mathbbm{1}-M$ in the effect space to ensure normalization. For the second task, the idea is to fix a large number of effects that are distributed in all directions on the Bloch sphere. We choose the following parametrization:
@@ 2.1 The impact of the number of states and measurements
Comparing the plots for pure and mixed states, it is immediately apparent that scenarios with pure states and projective measurements require fewer preparations and measurements to achieve high typicality of contextuality and also reach the saturation of $t_{(n,m,d;N)}\approx 100\%$ faster. For example, for pure states, we observe $t_{(n,m,d;N)}>99\%$ already for $(n=7,m=8)$, while for mixed states it only attains this threshold around $(n=18,m=16)$. Moreover, notice that if one only has access to $m=4$ measurements, typicality is $t_{(n,m,d;N)}>99\%$ for all $n>10$ pure preparations. Such saturation does not appear for mixed states in the range of $n$ and $m$ that we consider here. Additionally, the typicality of contextuality never reaches $t_{(n,m,d;N)}\approx 100\%$ for the mixed case in this range of numbers of states and measurements, although we do expect that it would if $n$ and $m$ were increased further. This numerical exploration, which pertains to Questions 1 and 2 in the Introduction, can be summarized as follows.
@@
For $(n,m,d=2;N=10^{6})$, typicality of contextuality is zero when $n=4$ for all $m$, for both pure and mixed preparations and measurements. When $n>4$, typicality of contextuality is already non-zero for pure states and quickly reaches $t_{(n,m,d;N)}\approx 100\%$ for $m\geq 4$. For mixed states and effects, a similar behavior is observed, but $t_{(n,m,d;N)}\approx 100\%$ is approached at a slower pace (i.e., larger values of $n$ and $m$ are needed).
In order to further explore Questions 1 and 2, we now move on to study how typicality of contextuality assessments change with the number of states when the number of measurements is large and chosen to approximate the full set of projective effects. This would provide a numerical visualization of the result in Ref. zhang2025reassessing , which states that if one has access to all possible measurements for a quantum system, linear dependence among a set of states is necessary (and in the case of pure states, sufficient [^5] For mixed states, linear dependence is not sufficient to imply contextuality. For example, an $\varepsilon$-ball of states around the maximally mixed state in the qubit space contains as many linear dependences as the unit ball but will nonetheless, for sufficiently small but non-zero $\varepsilon$, admit a noncontextual explanation in any prepare-and-measure scenario (since any contextual prepare-and-measure scenario becomes noncontextual under a finite amount of depolarizing noise schmid2018contextual ; marvian2020 ; ravi2015 ).) to generate contextuality.
@@
Examples of $n=5$ (a) pure states and (b) mixed states and their convex hulls (violet), measured by the same 92 projective measurements provided by Eq. (6) and their convex hull (blue, excluding the null and the unit effects). We therefore parametrize projectors of a qubit according to Eq. (6) and test the typicality for $(n,m=92,d=2;N=10^{6})$ for both pure and mixed states, with the caveat that now the measurements remain fixed for all trials. The results are displayed in Fig. 3. As expected, for $n=4$ we observe $t_{(n,m,d;N)}=0\%$ for both pure and mixed preparations, as generating contextuality in this case is only possible in very specific setups that will never happen in our random sampling. For $n\geq 5$, we have non-zero typicality of contextuality, and, in fact, for pure states, we observe $t_{(n,m,d;N)}\approx 100\%$ already for $n\geq 6$. The case with pure states and $n=5$ registers $t_{(n=5,m=92,d=2;N=10^{6})}=99.76\%$, and the calculated Wilson score interval gives us that the true value of $t_{(n,m,d)}$ lies above $98.99\%$ with 99% confidence. As mentioned before, for pure states, if one has access to all possible measurements, linear dependence is both necessary and sufficient to generate contextuality. Here, however, we do not observe a typicality of contextuality $t_{(n=5,m,d;N)}\approx 100\%$, even though five randomly sampled pure qubit states will always be linearly dependent. One could argue that this is due to the fact that, although large, the number of measurements in the experiment is finite. However, whether or not the possible associated contextuality is witnessed by the linear program in Ref. selby2024linear will depend on how dense the measurements are and on the numerical threshold used for the robustness. Hence, the fact that we have a finite number of measurements is not necessarily the only cause for typicality not being approximately 100%.
When sampling from mixed states, however, one requires an increasingly large number of states to reach $t_{(n,m,d;N)}\approx 100\%$, and $n=5$ displays a rather low typicality value. A visual comparison of one sampling for the pure and mixed cases for $n=5$ is given in Fig. 4: there we present (i) the convex hull of the measurements as a blue polytope, which is the same for both the pure state case and the mixed state case, and (ii) the convex hull of the $n=5$ sampled states as a purple polytope, when these are sampled from our methods for pure (Fig. 4(a)) and mixed states (Fig. 4(b)). It then appears that for the mixed case, the volume of the states polytope is considerably smaller than for the pure case. Therefore, what might be happening is that the polytope of mixed states can still be easily embedded into a simplex. For the mixed preparations, typicality reaches $t_{(n,m,d;N)}\geq 99\%$ for $n=10$, and $t_{(n,m,d;N)}\approx 100\%$ is observed starting at $n=14$. This reflects the fact that linear dependence is not sufficient to guarantee contextuality if one’s states are mixed.
@@
A useful way of interpreting these results is given by the recent work in Ref. zhang2025reassessing ; zhang2025quantifiers . While generalized (non)contextuality is usually viewed as a property to be assessed for a set of observed data or a complete experiment, Refs. zhang2025reassessing ; zhang2025quantifiers showed how this idea can be lifted to a notion of nonclassicality that applies even to a single component or fragment of an experiment (such as a single quantum state, a single quantum channel, a set of states, and so on). For example, a set of quantum states is classical if, when paired with all possible quantum measurements, it yields statistics that admit a generalized-noncontextual model. In the result just above, we have considered (a finite approximation of) the entire effect space of quantum theory, and so our result can be interpreted as approximately computing the typicality of nonclassicality for sets of quantum states (as opposed to of entire prepare-measure experiments).
@@ 2.2 The impact of purity and sharpness
We explore this by analyzing scenarios with the fixed number of measurements $m=20$. From the results in the previous section, we know that for any $m\geq 12$, typicality of contextuality reaches $t_{(n,m,d;N)}>99\%$ for $n\geq 7$ pure states. Moreover, for $m\geq 13$, typicality $t_{(n,m,d;N)}>99\%$ is already achieved for $n=6$. Our current question is: for $m=20$ measurements, how does the minimum number of states required to observe $t_{(n,m,d;N)}>99\%$ depend on the purity of the states and the set of allowed measurements? To address this, we examine different lower bounds on purity (measured by $\mathrm{tr}\left\{\rho^{2}\right\}\in\left[\frac{1}{d},1\right]$, where $\rho$ is the state of the system), thereby restricting how mixed the states can be. We study two cases: (i) restricting the purity of states while keeping the measurements projective and (ii) restricting the purity of both preparations and measurements. The results for $(n,m=20,d=2;N=10^{6})$ are presented in Fig. 5 for three lower bounds on purity: 0.5 $\frac{1}{d}$, 0.7, and 0.9, and two upper bounds on purity: 0.95 and 1.
@@
Figure 5: Minimal number of preparations $n$ needed to observe typicality $t_{(n,m=20,d=2;N=10^{6})}>99\%$ for $m=20$ measurements as a function of the lower bound on purity of the states with (a) sharp measurements and (b) unsharp measurements, for two different upper bounds on their purity. In (b), the bounds on sharpness are the same as the purity of the states.
For the case of projective measurements (Fig. 5(a)), one requires a considerably lower number of random states (compared to the case of POVMs in Fig. 5(b)) to generate contextuality with typicality $t_{(n,m,d;N)}>99\%$. From Fig. 5(a), notice that eliminating a thin slice of highly faithful states (i.e., the ones for which purity is between $0.95$ and $1$) does not jeopardize this ubiquity of nonclassicality—at least when the measurements are sharp. In fact, if the states are pure enough (purity larger than $0.9$), it does not matter whether one has access to this highly faithful slice or not. This is extremely relevant for experimental implementations, where high purity states may be costly or impossible to implement Girolami2019 ; Guryanova2020 . Our results thereby show that for 20 projective measurements, $n=9$ preparations already suffice for near-certain typicality when purity is at least $0.7$. If one can ensure at least $0.9$ purity in states, one requires only $n=7$ preparations.
Naturally, these numbers may change if the randomly sampled measurements are unsharp. As shown in Fig. 5(b), the minimal number of random preparations required for $t_{(n,m,d;N)}>99\%$ with 20 random POVMs is as high as $n=16$ for the case where no lower or upper bounds on purity are imposed. Notice that because they are a measure-zero subset of the full effect space, projective effects are never sampled in this case. It is therefore interesting to notice that the minimal number of preparations is $n=7$ for states and effects with purity between $0.9$ and $1$. This is the same number as the one observed for projective measurements and states in this slice, suggesting that, within this range of purity, contextuality is easy to generate even without access to projective measurements.
@@
All the values for the minimal number of states depend on the number of fixed measurements $m$, as well as the other parameters involved in our computations and discussed in detail in Appendix C. In order to facilitate the investigation for particular scenarios, we provide the open-source Python implementation that generated all of our plots in Ref. github , along with a function that can be used to obtain the minimal number of preparations to be sampled for $t_{(n,m,d;N)}>99\%$ given $m$, $d$, NN and upper and lower bounds for the purity of states (and possibly effects). A summary of this repository is provided in Appendix D.

### v2

@@ 1.2 Methods
To study the typicality of quantum contextuality, we therefore perform extensive simulations of contextuality certification. We characterize each simulation by a tuple of parameters $(n,m,d;N)$, where $n$ is the number of preparations and $m$ is the number of binary measurements that comprise the prepare-and-measure scenario, $d$ is the dimension of the Hilbert space associated to the quantum system being investigated, and $N$ is the number of sampling iterations and assessments of contextuality. We begin by randomly sampling $n$ density matrices of dimension $d\times d$ and $m$ pairs of POVM elements, each of which rescaled to ensure normalization (amounting to a total of $2m$ effects). We then employ the linear program selby2024linear to determine whether the statistics specified by $(n,m,d)$ admit a generalized-noncontextual model.
@@
Notice that, throughout this work, stating $t_{(n,m,d;N)}\approx 100\%$ does not mean that contextuality will always be found for $n$ preparations and $m$ binary measurements for a quantum system of dimension $d$, but rather that out of $N$ tries, every single one of them exhibited contextuality by our certification criteria. To calculate the confidence level of the true value of $t_{(n,m,d)}$, we use the Wilson score interval wilson , which we explain in detail in Appendix C. In summary, when we say $t_{(n,m,d;N=10^{6})}\approx 100\%$ in this work, this criterion ensures that the typicality of contextuality $t_{(n,m,d)}$ is at least 99.999% with 99% level of confidence. In this paper, we focus on qubit states $(d=2)$, but the general approach and the simulation toolbox that we share in an online repository are also suitable for higher dimensions.
To generate random states, we employ the qutip toolbox lambert2024qutip5quantumtoolbox . Pure random states distributed with a Haar measure are generated by applying a Haar random unitary to a fixed pure state (rand_ket function). To generate mixed states, we follow the algorithm designed in Ref. Bruzda2009 that generates Ginibre random density operators of fixed rank (rand_dm function). Since only full-rank density matrices have nonzero measure Bengtsson2017 , sampling from mixed states will never lead to pure states. More details on random sampling are provided in Appendix B. To sample the effects, there are two approaches that we employ depending on the task— randomly sampling effects or fixing a large number of (somewhat uniform) effects. For the first approach, we employ qutip’s random unitary generator to sample projective measurements by rotating an orthornomal basis; and numpy’s random generator function to sample POVMs by generating sets of Hermitian operators from the Ginibre ensamble, then enforcing measurement normalization kukulski2021 (See Appendix B). For the second task, the idea is to fix a large number of effects that are distributed in all directions on the Bloch sphere. We choose the following parametrization:
@@ 2.1 The impact of the number of states and measurements
Comparing the plots for pure and mixed states, it is immediately apparent that scenarios with pure states and projective measurements require fewer preparations and measurements to achieve high typicality of contextuality and also reach the saturation of $t_{(n,m,d;N)}\approx 100\%$ faster. For example, for pure states and effects, we observe $t_{(n,m,d;N)}>99\%$ already for $(n=7,m=8)$, while for mixed states and effects values above this range are never achieved for the considered values of $n$ and $m$, and the maximum value registered for this case is $t_{(n,m,d;N)}=97.7\%$. Moreover, notice that if one only has access to $m=4$ measurements, typicality is $t_{(n,m,d;N)}>99\%$ for all $n>10$ pure preparations. Additionally, contrarily to the pure case, the typicality of contextuality never reaches $t_{(n,m,d;N)}\approx 100\%$ for the mixed case in this range of numbers of states and measurements, although we do expect that it would if $n$ and $m$ were increased further. This numerical exploration, which pertains to Questions 1 and 2 in the Introduction, can be summarized as follows.
@@
For $(n,m,d=2;N=10^{6})$, typicality of contextuality is zero when $n=4$ for all $m$, for both (i) pure states and sharp measurements and (ii) mixed states and unsharp measurements. When $n>4$, typicality of contextuality is already non-zero for pure states and sharp measurements, and quickly reaches $t_{(n,m,d;N)}\approx 100\%$ for $m\geq 4$. For mixed states and POVMs, a similar behavior is observed, but $t_{(n,m,d;N)}\approx 100\%$ is approached at a slower pace (i.e., larger values of $n$ and $m$ are needed).
In order to further explore Questions 1 and 2, we now move on to study how typicality of contextuality assessments change with the number of states when the number of measurements is large and chosen to approximate the full set of projective effects. This would provide a numerical visualization of the result in Ref. zhang2025reassessing , which states that if one has access to all possible measurements for a quantum system, linear dependence among a set of states is necessary (and in the case of pure states, sufficient [^3] For mixed states, linear dependence is not sufficient to imply contextuality. For example, an $\varepsilon$-ball of states around the maximally mixed state in the qubit space contains as many linear dependences as the unit ball but will nonetheless, for sufficiently small but non-zero $\varepsilon$, admit a noncontextual explanation in any prepare-and-measure scenario (since any contextual prepare-and-measure scenario becomes noncontextual under a finite amount of depolarizing noise schmid2018contextual ; marvian2020 ; ravi2015 ).) to generate contextuality. We start the discussion with an analytical result:
Proposition 2.
For any finite $(n,m,d)$ with $n$ preparations and $m$ measurements, $t_{(n,m,d)}<1$. This is holds even when sampling only pure states and effects.
A proof is given in Appendix D.
This reiterates that assessments of the form $t_{(n,m,d;N)}\approx 100\%$ throughout this work do not translate to the theoretical typicality $t_{(n,m,d)}=1$. As we see now, however, with a large but finite number of measurements one consistently gets $t_{(n,m,d;N)}\approx 100\%$. Later on, in Prop. 3, we will show that $\lim_{m\to\infty}t_{(n,m,d=2)}=1$ for $n>4$.
@@
We parametrize projectors of a qubit according to Eq. (6) and test the typicality for $(n,m=92,d=2;N=10^{6})$ for both pure and mixed states, with the caveat that now the measurements remain fixed for all trials. The results are displayed in Fig. 3. As expected, for $n=4$ we observe $t_{(n,m,d;N)}=0\%$ for both pure and mixed preparations, as generating contextuality in this case is only possible in very specific setups that will never happen in our random sampling. For $n\geq 5$, we have non-zero typicality of contextuality, and, in fact, for pure states, we observe $t_{(n,m,d;N)}\approx 100\%$ already for $n\geq 6$. The case with pure states and $n=5$ registers $t_{(n=5,m=92,d=2;N=10^{6})}=99.76\%$, and the calculated Wilson score interval gives us that the true value of $t_{(n,m,d)}$ lies above $98.99\%$ with 99% confidence. As mentioned before, for pure states, if one has access to all possible measurements, linear dependence is both necessary and sufficient to generate contextuality. Here, however, we do not observe a typicality of contextuality $t_{(n=5,m,d;N)}\approx 100\%$, even though five randomly sampled pure qubit states will always be linearly dependent. One could argue that this is due to the fact that, although large, the number of measurements in the experiment is finite. However, whether or not the possible associated contextuality is witnessed by the linear program in Ref. selby2024linear will depend on how dense the measurements are and on the numerical threshold used for the robustness. Hence, the fact that we have a finite number of measurements is not necessarily the only cause for typicality not being approximately 100%.
When sampling from mixed states, however, one requires an increasingly large number of states to reach $t_{(n,m,d;N)}\approx 100\%$, and $n=5$ displays a rather low typicality value. A visual comparison of one sampling for the pure and mixed cases for $n=5$ is given in Fig. 4: there we present (i) the convex hull of the measurements as a blue polytope, which is the same for both the pure state case and the mixed state case, and (ii) the convex hull of the $n=5$ sampled states as a purple polytope, when these are sampled from our methods for pure (Fig. 4(a)) and mixed states (Fig. 4(b)). It then appears that for the mixed case, the volume of the states polytope is considerably smaller than for the pure case. Therefore, a reasonable explanation is that the polytope of mixed states can still be easily embedded into a simplex. For the mixed preparations, typicality reaches $t_{(n,m,d;N)}\geq 99\%$ for $n=10$, and $t_{(n,m,d;N)}\approx 100\%$ is observed starting at $n=14$. This reflects the fact that linear dependence is not sufficient to guarantee contextuality if one’s states are mixed.
@@
A useful way of interpreting these results is given by the recent work in Ref. zhang2025reassessing ; zhang2025quantifiers . While generalized (non)contextuality is usually viewed as a property to be assessed for a set of observed data or a complete experiment, Refs. zhang2025reassessing ; zhang2025quantifiers showed how this idea can be lifted to a notion of nonclassicality that applies even to a single component or fragment of an experiment (such as a single quantum state, a single quantum channel, a set of states, and so on). For example, a set of quantum states is classical if, when paired with all possible quantum measurements, it yields statistics that admit a generalized-noncontextual model. In the result just above, we have considered (a finite approximation of) the entire effect space of quantum theory, and so our result can be interpreted as approximately computing the typicality of nonclassicality for sets of quantum states (as opposed to of entire prepare-measure experiments). In fact, we now prove that in the asymptotic limit, the analytical typicality reaches $\lim_{m\to\infty}t_{(n,m,d)}=1$.
Proposition 3.
Consider samples of $n$ pure states and $m$ pure effects of a $d$-dimensional quantum system. For any $n>d^{2}$ number of randomly sampled pure states,
$\lim_{m\to\infty}t_{(n,m,d)}=1.$ (7)
Moreover, for $m>d^{2}$ randomly sampled pure effects,
$\lim_{n\to\infty}t_{(n,m,d)}=1.$ (8)
The proof is given in Appendix D.
The case investigated in this work of sampling dichotomic measurements naturally applies to this proof (in which case $m\geq d^{2}/2$ already ensures that the scenario is not ruled out by Lemma 1). A useful property stemming from this result is that $t_{(n,m,d)}$ grows monotonically with $n$ and $m$.
Corollary 4.
For any $n>d^{2}$ number of randomly sampled pure states and any real number $t<1$, there is a number $m\in\mathbb{N}$ of randomly sampled pure effects such that $t_{(n,m,d)}>t$. Similarly, for any $m>d^{2}$ randomly sampled pure effects, there is a number $n\in\mathbb{N}$ of randomly sampled pure states such that $t_{(n,m,d)}>t$.
@@ 2.2 The impact of purity and sharpness
We explore this by analyzing scenarios with the fixed number of measurements $m=20$. From the results in the previous section, we know that for any $m\geq 12$, typicality of contextuality reaches $t_{(n,m,d;N)}>99\%$ for $n\geq 7$ pure states. Moreover, for $m\geq 13$, typicality $t_{(n,m,d;N)}>99\%$ is already achieved for $n=6$. Our current question is: for $m=20$ measurements, how does the minimum number of states required to observe $t_{(n,m,d;N)}>99\%$ depend on the purity of the states and the set of allowed measurements? To address this, we examine different lower bounds on purity (measured by $\mathrm{tr}\left\{\rho^{2}\right\}\in\left[\frac{1}{d},1\right]$, where $\rho$ is the state of the system), thereby restricting how mixed the states can be. We study two cases: (i) restricting the purity of states while keeping the measurements projective and (ii) restricting the purity of both states and the sharpness of measurements. The results for $(n,m=20,d=2;N=10^{6})$ are presented in Fig. 5 for three lower bounds on purity/sharpness: 0.5 $\frac{1}{d}$, 0.7, and 0.9, and two upper bounds on purity/sharpness: 0.95 and 1.
@@
Figure 5: Minimal number of states $n$ needed to observe typicality $t_{(n,m=20,d=2;N=10^{6})}>99\%$ for $m=20$ measurements as a function of the lower bound on purity of the states with (a) sharp measurements and (b) unsharp measurements, for two different upper bounds on their purity. In (b), the bounds on sharpness are the same as the purity of the states.
For the case of projective measurements (Fig. 5(a)), one requires a considerably lower number of random states (compared to the case of POVMs in Fig. 5(b)) to generate contextuality with typicality $t_{(n,m,d;N)}>99\%$. From Fig. 5(a), notice that eliminating a thin slice of highly faithful states (i.e., the ones for which purity is between $0.95$ and $1$) does not jeopardize this ubiquity of nonclassicality—at least when the measurements are sharp. In fact, if the states are pure enough (purity larger than $0.9$), it does not matter whether one has access to this highly faithful slice or not. Our results thereby show that for 20 projective measurements, $n=9$ preparations already suffice for near-certain typicality when purity is at least $0.7$. If one can ensure at least $0.9$ purity in states, one requires only $n=7$ preparations.
Naturally, these numbers may change if the randomly sampled measurements are unsharp. As shown in Fig. 5(b), the minimal number of random preparations required for $t_{(n,m,d;N)}>99\%$ with 20 random POVMs is as high as $n=25$ for the case where no lower or upper bounds on purity are imposed. Notice that because they are a measure-zero subset of the full effect space, projective effects are never sampled in this case. It is therefore interesting to notice that the minimal number of preparations is $n=7$ for states and effects with purity between $0.9$ and $1$. This is the same number as the one observed for projective measurements and states in this slice, suggesting that, within this range of purity, contextuality is easy to generate even without access to projective measurements. This is extremely relevant for experimental implementations, where high purity states or sharp measurements may be costly or impossible to implement Girolami2019 ; Guryanova2020 .
@@
All the values for the minimal number of states depend on the number of fixed measurements $m$, as well as the other parameters involved in our computations and discussed in detail in Appendix C. In order to facilitate the investigation for particular scenarios, we provide the open-source Python implementation that generated all of our plots in Ref. github , along with a function that can be used to obtain the minimal number of preparations to be sampled for $t_{(n,m,d;N)}>99\%$ given $m$, $d$, NN and upper and lower bounds for the purity of states (and possibly effects). A summary of this repository is provided in Appendix E.